OpenStudy (anonymous):

why is the sum of the two numbers -1/5 and .20 0 ? i dont get it

OpenStudy (anonymous):

Because \(\dfrac{1}{5}=0.20\). So, if you have \(-\dfrac{1}{5}+0.20=-\dfrac{1}{5}+\dfrac{1}{5}=0\).

OpenStudy (anonymous):

so they all cancel out?

OpenStudy (anonymous):

Yes. Do you understand why \(\dfrac{1}{5}=0.20\)?

OpenStudy (anonymous):

yes because .20 is 1/5 of a hundred